# Efficient Bayesian Additive Regression Models For Microbiome Studies

### v1

@@ 3.4 Inferring Model Hyperparameters
Hyperparameter selection remains an outstanding problem in MLTPs (Silverman et al., 2022). Here we demonstrate how the same Laplace approximation used in the Collapse step of the CU sampler can be used to approximate marginal likelihoods and thereby estimate hyperparameters via Maximum Marginal Likelihood (MML) or penalized MML estimation.
Let $\Omega$ denote a set of hyperparameters in the joint MLTP distribution: $p(\mathbf{H},\Phi,\mathbf{Y}\mid\Omega)$. These can be parameters in a kernel, a mean function, or even in the log-ratio transformation $\phi$. The MML estimate for $\Omega$ is
@@
where $\hat{\mathbf{H}}_{\Omega}$ denotes the MAP estimate for $\mathbf{H}$ evaluated in the observed set $\{1,\dots,N\}$ which depends on the hyperparameters $\Omega$. From a computational standpoint, this Laplace approximation to the marginal likelihood is essentially free given that we are already computing the MAP estimate and a factorization of the Hessian matrix as part of the CU sampler (Silverman et al., 2022).
While the MML estimate already provides a degree of penalization (Wood, 2011), it can be helpful to add an additional penalty term to the marginal likelihood when defining weakly identified MultiAddGPs. As we demonstrate in Section 5 and Supplementary Section 5, we often choose this penalty term as the log density of a prior for the hyperparameters.
In practice, we find that the CU sampler is efficient enough to make optimizing Equation (7) (with or without added penalization) practical. While we have found that various optimization routines can be used to perform the optimization, the problem is often non-convex, especially when estimating bandwidth parameters in kernel functions. For this reason, we generally recommend Bayesian optimization procedures which can optimize non-convex functions while simultaneously reducing the number of function evaluations compared to more typical optimizers (e.g., the L-BFGS optimizer). We have tested and found Bayesian optimization procedures generally work well when $\Omega$ is of low to moderate dimension (e.g., $\leq 10$). When $\Omega$ is higher dimensional, alternative procedures may be needed. In this study, Bayesian Optimization was performed using the ParBayesianOptimization package with the acquisition function set to the upper confidence bound, starting with 10 initial points and proceeding through 20 iterations with all other parameters set to their default values.
@@ 3.5 Software Availability
We extend the R package fido (Silverman et al. (2022)) to efficiently infer MultiAddGPs models. The fido package implements the extended CU sampler described above using optimized C++ code. MAP estimation of $\hat{H}$ is performed using the L-BFGS optimizer. Hyperparameter selection by MML is performed by maximizing the marginal likelihood using Bayesian optimization implemented in the ParBayesianOptimization package. All code required to reproduce the results of the this article is available at https://github.com/Silverman-Lab/MultiAddGPs.
@@ 4 Simulations
We simulated a suite of longitudinal studies of microbiota with varying numbers of taxa $D\in\{3,\dots,100\}$ and samples $N\in\{20,\dots,1000\}$. We simulated microbial composition influenced by batch effects, daily periodicity (e.g., circadian rhythm; Heddes et al. (2022)), and longer-term trends. Full simulation details are provided in Supplementary Section 4.
@@
In Figure 1 we show a small simulation $D=4$ and $N=600$ for ease of visualization. We use $t_{n}$ to denote the time at which sample $n$ was obtained. For inference, we specify a MultiAddGP model $\mathbf{F}_{n}=\mathbf{b}_{0}+\mathbf{b}_{1}x^{(\text{batch})}_{n}+\mathbf{f}^{(\text{periodic})}(t_{n})+\mathbf{f}^{(\text{trend})}(t_{n})$ as follows. For covariates, we set $\mathbf{X}_{\cdot n}=[1\;x^{(\text{batch})}_{n}]^{T}$, $\mathbf{Z}^{(\text{periodic})}_{\cdot n}=t_{n}$, and $\mathbf{Z}^{(\text{trend})}_{\cdot n}=t_{n}$. For priors we set $\mathbf{B}=[\mathbf{b}_{0}=2.7;\mathbf{b_{1}}=1]$. Both $\mathbf{f}^{(\text{periodic})}$ and $\mathbf{f}^{(\text{trend})}$ were given matrix-normal process priors with mean function $\boldsymbol{\Theta}^{(k)}=\mathbf{0}$. A periodic kernel $K_{\text{period}}(t,t^{\prime})=\sigma_{\text{period}}\exp\left(-\frac{2\sin^{2}\left(\frac{\pi|t-t^{\prime}|}{p}\right)}{\rho_{\text{period}}^{2}}\right)$ was used in the prior for $\mathbf{f}^{(\text{periodic})}$ and a squared exponential $K_{\text{trend}}(t,t^{\prime})=\sigma_{\text{trend}}^{2}\exp(-\frac{(t-t^{\prime})^{2}}{2\rho_{\text{trend}}^{2}})$ was used for $\mathbf{f}^{(\text{trend})}$. Hyperparameters $\Omega=\{\sigma_{\text{period}},\rho_{\text{period}},p,\sigma_{\text{trend}},\rho_{\text{trend}}\}$ were selected using MML estimation.
For comparison, we created a nearly identical model that ignored uncertainty due to counting and assumed the data was transformed Gaussian. We implemented this model by setting $\mathbf{H}_{\cdot n}=\phi(Y_{\cdot n}+0.5)$ and proceeding with the uncollapse step of MultiAddGPs directly; skipping sampling the posterior of the collapsed form. We call this model the Normal Additive GP (NAddGP) model.
@@
We compared posterior estimates from the MultiAddGP and NAddGP models to emphasize the importance of modeling uncertainty due to counting. Figure 1 shows that the MultiAddGP almost perfectly recovered the true decomposition whereas the NAddGP substantially underestimated the amplitude of the periodic component and long-term trend.
@@
Figure 1: MultiAddGPs successfully decompose simulated microbiome time-series. The NAddGP model is identical to the MultiAddGP model but ignores uncertainty due to counting by modeling the observed data as transformed Gaussian. Panels A, B, and C represent individual decomposed components associated with each covariate. Panel D illustrates the cumulative effect of all components.
@@
Figure 2 shows these findings generalize as $N$ and $D$ increase. As the posterior of these high-dimensional models cannot be easily visualized, we quantified model performance based on the coverage of posterior 95% intervals with respect to the true function decomposition. Since both MultiAddGP and NAddGP are Bayesian models we do not expect that these intervals will cover the truth with 95% probability. As a result, we focus on the ratio of coverage between the MultiAddGP and the NAddGP. Positive values of this coverage ratio indicate that the MultiAddGP model covers the truth more often than the NAddGP model. In all simulations, at all sample sizes $N$ and number of taxa $D$, the MultiAddGP models covered the truth more frequently than the NAddGP models.
@@
Figure 2: At all tested dimensions $N$ and $D$, posterior intervals from MultiAddGPs cover the truth more frequently than NAddGPs. The first row illustrates how data sparsity varied with dimensions $N$ and $D$ in our simulation studies. The second row shows the ratio between coverage of 95% Credible intervals from MultiAddGPs compared to NAddGPs. Each datapoint represents the mean over three simulations. The ratio is always positive illustrating MultiAddGPs cover the truth substantially more than NAddGPs. Coverage ratios for each of the decomposed components Batch, Hourly, and Daily and the cumulative function $F$ are shown.
5 Analysis of Artificial Human Guts
@@
We used MultiAddGPs to reanalyze on a previously published longitudinal study of four artificial gut models (Silverman et al., 2018). This study consists of over 500 samples from $4$ artificial gut vessels irregularly spaced through time. The key feature of this study was the starvation of Vessels 1 and 2 that occurred over days 11 to 13 while Vessels 3 and 4 received no intervention and therefore act as controls. We have previously modeled these data using Multinomial Logistic-Normal models (Silverman et al., 2022, 2018) and demonstrated that microbiota demonstrate a prolonged recovery phase after feeding resumes. However, prior models were limited and could not separate long-term trend from the effect of starvation. We use a MultiAddGP model to overcome this limitation and provide novel insights into these microbial communities.
@@
Data preprocessing followed the analysis of Silverman et al. (2018) and resulted in 10 bacterial families and 537 samples. We modeled these data as consisting of two overlapping temporal processes, a vessel-specific long-term trend $f^{(\text{base},v)}$ and a vessel-specific function representing the effect of starvation $f^{(\text{disrupt},v)}$. We used the following MultiAddGP model:
@@
where $\boldsymbol{\Xi}$ was chosen to reflect a prior belief that more evolutionary similar taxa would behave more similarly (Silverman et al., 2017), $\zeta=D+5$ reflects weak prior knowledge in this assumption, and $\odot$ represents elementwise multiplication of kernels. The kernel $\boldsymbol{\Gamma}^{(v)}$ was block diagonal with all non-zero elements equal to 1: it was used to model conditional independence between the vessels. $\boldsymbol{\Gamma}^{(\text{base})}$ was modeled as a squared exponential kernel to capture long-term non-linear trends. In contrast, $\boldsymbol{\Gamma}^{(\text{distrup})}$ was modeled using a rational quadratic kernel which was forced to zero before day 11 to reflect our knowledge that starvation started on day 11. Hyperparameters $\Omega=\{\sigma_{\text{base}},\rho_{\text{base}},\sigma_{\text{disrupt}},\rho_{\text{disrupt}}\}$ were selected using penalized MML. Full details on penalization, kernel, and mean functions are provided in Supplementary Section 5.
@@
Posterior estimates for the cumulative function $F$ (long-term trend plus starvation effect) are shown in Figure 3. Posterior estimates for the decomposed starvation effect are shown in Figure 4. Consistent with prior reports, the Rikenellaceae family demonstrates a substantial decrease during the initial starvation period followed by a slow recovery to baseline (Silverman et al., 2018; Sun and Zhou, 2024). This finding is notable given that decreases in Rikenellaceae are also observed in obese individuals (Okeke, Roland and Mullin, 2014; Peters et al., 2018; Tavella et al., 2021). As obesity and starvation seem like opposite conditions, decreased Rikenellaceae relative abundance in both conditions suggest that host factors not present in these artificial gut systems but present in in vivo obesity studies (e.g., the host immune system) may be driving the decreases in obesity.
@@
Our MultiAddGPs also reveal new features of this data and suggest novel biology. For example, our model finds strong evidence that the Rikenellaceae relative abundance in the community decreases drastically during the initial starvation event followed by a later over-correction where their abundance ultimately increases before trending back to pre-starvation levels. These types of over-corrections are predicted by ecological predator-prey models which can lead to damped oscillations in response to exogenous stimuli (Samuelson, 1971). Still, to best of our knowledge, this is the first study to demonstrate phenomena consistent with that theory within in vitro systems.
@@
Figure 3: MultiAddGP Posterior for the Smoothed State $\mathbf{F}$ of an Artificial Gut Study. As described in the main text, $\mathbf{F}$ was modeled as the additive combination of a slowly-varying temporal trend and a more variable yet short-lived starvation effect in Vessels 1 and 2 which started on Day 11. Posterior mean and 95% credible intervals of $\mathbf{F}$ are shown for each of the four artificial gut vessels. We plot these results with respect to the Centered Log-Ratio (CLR) Coordinates of each bacterial family. The starvation event, which occurred between Days 11 and 13 in Vessels 1 and 2, is highlighted in Gray.
@@
Figure 4: MultiAddGPs Posterior for the Starvation effect ($\mathbf{f}^{(disrupt)}$) in Artificial Gut Study. Posterior mean and 95% credible intervals are shown for all taxa.

### v2

@@ 2.3 Model Identification
We find this method is computationally efficient, easy to implement, and provides similar results to alternative methods.
@@ 2.4 Inferring Model Hyperparameters
Hyperparameter selection remains an outstanding problem in MLTPs (Silverman et al. 2022). Here we demonstrate how the same Laplace approximation used in the Collapse step of the CU sampler can be used to approximate marginal likelihoods and thereby estimate hyperparameters via Maximum Marginal Likelihood (MML) or penalized MML estimation.
Let $\Omega$ denote a set of hyperparameters in the joint MLTPs distribution: $p(\mathbf{H},\Phi,\mathbf{Y}\mid\Omega)$. These can be parameters in a kernel, a mean function, or even in the log-ratio transformation $\phi$. The MML estimate for $\Omega$ is
@@
where $\hat{\mathbf{H}}_{\Omega}$ denotes the MAP estimate for $\mathbf{H}$ evaluated in the observed set $\{1,\dots,N\}$ which depends on the hyperparameters $\Omega$. From a computational standpoint, this Laplace approximation to the marginal likelihood is essentially free given that we are already computing the MAP estimate and a factorization of the Hessian matrix during the CU sampler. Beyond MML estimation, it is simple to add additional penalization when defining weakly identified MultiAddGPs.
In practice, we estimate hyperparameters by maximizing Equation (6), either with or without additional penalization. While various optimization routines can be used to perform the optimization, the problem is often non-convex, especially when estimating bandwidth parameters in kernel functions. For this reason, we generally recommend Bayesian optimization procedures, which can optimize non-convex functions while simultaneously reducing the number of function evaluations compared to more typical optimizers (e.g., the L-BFGS optimizer). We have tested and found Bayesian optimization procedures generally work well when $\Omega$ is of low to moderate dimension (e.g., $\leq 10$). When $\Omega$ is higher dimensional, alternative procedures may be needed. In this study, Bayesian Optimization was performed using the ParBayesianOptimization R package with the acquisition function set to the upper confidence bound, starting with 10 initial points and proceeding through 20 iterations with all other parameters set to their default values.
@@ 2.5 Software Availability
We have made MultiAddGPs available through the R package fido (Silverman & Nixon 2021), where it can be accessed via the basset function. The fido package implements the extended CU sampler described above using optimized C++ code. MAP estimation of $\hat{H}$ is performed using the L-BFGS optimizer. Hyperparameter selection by MML is performed by maximizing the marginal likelihood using Bayesian optimization implemented in the ParBayesianOptimization R package (Wilson 2022). All code required to reproduce the results of this article is available at https://github.com/Silverman-Lab/MultiAddGPs.
@@ 3 Simulations
To evaluate the performance of the proposed model relative to existing methods, we conduct two simulation studies. The first study highlights the importance of accounting for uncertainty introduced by the measurement process, which is ignored by LongGP (Cheng et al. 2019) . The second study compares the performance of MultiAddGPs to Lgpr (Timonen et al. 2021). However, due to the high computational cost of Lgpr, we were unable to scale it to the full simulation setting. Instead, we perform a smaller-scale simulation to compare MultiAddGPs and Lgpr. Full simulation details are provided in Supplementary Section 5.
3.1 Simulation 1
We simulated longitudinal microbiome studies with varying numbers of taxa $D\in\{3,\dots,100\}$ and samples $N\in\{50,\dots,1000\}$. We simulated microbial composition influenced by batch effects, daily periodicity (e.g., circadian rhythm; Heddes et al. (2022)), and longer-term trends. For this simulation study we did not compare to Lgpr or LongGP for two reasons: (1) it was too computationally intensive and could not scale to the larger simulation sizes; (2) it could not model periodic components as it lacked periodic kernels. Instead, we created a new method Normal Additive GPs (NAddGPs) which assumes the data are transformed Gaussian and are therefore reminiscent of LongGP. NAddGPs are nearly identical to MultiAddGPs but assume each $\mathbf{H}_{\cdot n}$ is observed and equal to $\mathbf{H}_{\cdot n}=\phi(\mathbf{Y}_{\cdot n}+0.5\cdot\mathbf{1}_{D})$ and therefore proceed directly to the uncollapse step, skipping the collapsed step of our CU sampler.
@@ 3.1 Simulation 1
In Figure 1 we show a small simulation $D=4$ and $N=600$ for ease of visualization. We use $t_{n}$ to denote the time at which sample $n$ was obtained. For inference, we specified MultiAddGPs and NAddGPs models as $\mathbf{F}_{n}=\mathbf{b}_{0}+\mathbf{b}_{1}x^{(\text{batch})}_{n}+\mathbf{f}^{(\text{periodic})}(t_{n})+\mathbf{f}^{(\text{trend})}(t_{n})$. For covariates, we set $\mathbf{X}_{\cdot n}=[1\;x^{(\text{batch})}_{n}]^{T}$, $\mathbf{Z}^{(\text{periodic})}_{\cdot n}=t_{n}$, and $\mathbf{Z}^{(\text{trend})}_{\cdot n}=t_{n}$. For priors we set $\mathbf{B}=[\mathbf{b}_{0}=2.7;\mathbf{b_{1}}=1]$. Both $\mathbf{f}^{(\text{periodic})}$ and $\mathbf{f}^{(\text{trend})}$ were given matrix-normal process priors with mean function $\boldsymbol{\Theta}^{(k)}=\mathbf{0}$. A periodic kernel $K_{\text{period}}(t,t^{\prime})=\sigma_{\text{period}}\exp\left(-\frac{1}{\rho_{\text{period}}^{2}}2\sin^{2}\left(\frac{\pi|t-t^{\prime}|}{p}\right)\right)$ was used in the prior for $\mathbf{f}^{(\text{periodic})}$ and a squared exponential $K_{\text{trend}}(t,t^{\prime})=\sigma_{\text{trend}}^{2}\exp(-\frac{1}{2\rho_{\text{trend}}^{2}}(t-t^{\prime})^{2})$ was used for $\mathbf{f}^{(\text{trend})}$. Hyperparameters $\Omega=\{\sigma_{\text{period}},\rho_{\text{period}},p,\sigma_{\text{trend}},\rho_{\text{trend}}\}$ were selected using MML estimation.
@@
We compared posterior estimates from the MultiAddGPs and NAddGPs models to emphasize the importance of modeling uncertainty due to counting. Figure 1 shows that the MultiAddGPs almost perfectly recovered the true decomposition whereas the NAddGPs substantially underestimated the amplitude of the periodic component and long-term trend.
@@
Figure 1: MultiAddGPs successfully decompose simulated microbiome time-series. The NAddGPs model is identical to the MultiAddGPs model but ignores uncertainty due to counting by modeling the observed data as transformed Gaussian. Panels A, B, and C represent individual decomposed components associated with each covariate. Panel D illustrates the cumulative effect of all components. In all panels, MultiAddGPs more accurately recover the true underlying signal compared to NAddGPs.
@@
Figure 2 shows these findings generalize as $N$ and $D$ increase. As the posterior of these high-dimensional models cannot be easily visualized, we quantified model performance based on the coverage of posterior 95% intervals with respect to the true function decomposition. Since both MultiAddGPs and NAddGPs are Bayesian models, we do not expect that these intervals will cover the truth with 95% probability. As a result, we focus on the ratio of coverage between the MultiAddGPs and the NAddGPs. Positive values of this coverage ratio indicate that the MultiAddGPs model covers the truth more often than the NAddGPs model. In all simulations, at all sample sizes $N$ and number of taxa $D$, the MultiAddGPs models covered the truth far more frequently than the NAddGPs models.
@@
Figure 2: At all tested dimensions $N$ and $D$, posterior intervals from MultiAddGPs cover the truth more frequently than NAddGPs. The first row illustrates how data sparsity varied with dimensions $N$ and $D$ in our simulation studies. The second row shows the ratio between coverage of 95% Credible intervals from MultiAddGPs compared to NAddGPs. Each datapoint represents the mean over three simulations. The ratio is always positive, illustrating that MultiAddGPs cover the truth substantially more than NAddGPs. Coverage ratios for each of the decomposed components Batch, Hourly, and Daily and the cumulative function $\mathbf{F}$ are shown.
3.2 Simulation 2
To compare with Lgpr, which allows for negative binomial likelihoods, we conducted a smaller simulation with $D=3$ taxa and $N=50$ samples. Since Lgpr employs predefined kernel structures and does not incorporate a periodic kernel, we designed the simulation based on an alternative latent function $\mathbf{F}_{n}=\mathbf{b_{0}}+\mathbf{b_{1}}x_{n}^{(\text{batch})}+\mathbf{f}^{(\text{non-stationary})}(t_{n})+\mathbf{f}^{(\text{trend})}(t_{n})$. We chose priors for Lgpr that matched the priors used for our MultiAddGPs.
Lgpr required approximately seven hours to analyze a single taxon on a high-performance computing cluster, resulting in an effective sample size of only 100 (Gelman et al. 2015). In contrast, MultiAddGPs simultaneously analyzed all taxa in about one minute and produced an effective sample size of 2000. In summary, our method produces effective samples at over 240,000 times the rate of Lgpr.
Our methods not only demonstrate increased efficiency, but the posterior estimates obtained using MultiAddGPs also align more closely with the true function decomposition compared to those derived from Lgpr (see Supplementary Figure 3).
4 Application to Artificial Human Guts
@@
We used MultiAddGPs to reanalyze a previously published longitudinal study of four artificial gut models (Silverman et al. 2018). This study consists of over 500 samples from $4$ artificial gut vessels irregularly spaced through time. The key feature of this study was the starvation of Vessels 1 and 2 that occurred over days 11 to 13 while Vessels 3 and 4 received no intervention and therefore act as controls. The study were previously be analyzed by Silverman et al. (2022) and Silverman et al. (2018), who show that microbiota demonstrate a prolonged recovery phase after feeding resumes. However, previous models could not disentangle long-term temporal trends from the specific effects of starvation, leaving the dynamics of the recovery phase poorly understood. We use a MultiAddGPs model to overcome this limitation and provide novel insights into these microbial communities.
@@
Data preprocessing followed the analysis of Silverman et al. (2018) and resulted in 10 bacterial families and 537 samples. We modeled these data as consisting of two overlapping temporal processes, a vessel-specific long-term trend $f^{(\text{base},v)}$ and a vessel-specific function representing the effect of starvation $f^{(\text{disrupt},v)}$. We used the following MultiAddGPs model:
@@
where $\boldsymbol{\Xi}$ was chosen following Silverman et al. (2022) to reflect the phylogenetic relationships between taxa. $\zeta=D+5$ was chosen to reflect weak prior knowledge of the importance of these relationships. Within the kernel functions, $\odot$ represents elementwise multiplication of kernels. The kernel $\boldsymbol{\Gamma}^{(v)}$ was block diagonal with all non-zero elements equal to 1: it was used to model conditional independence between the vessels. $\boldsymbol{\Gamma}^{(\text{base})}$ was modeled as a squared exponential kernel to capture long-term non-linear trends. In contrast, $\boldsymbol{\Gamma}^{(\text{distrup})}$ was modeled using a rational quadratic kernel, which was forced to zero before day 11 to reflect our knowledge that starvation started on day 11. Hyperparameters $\Omega=\{\sigma_{\text{base}},\rho_{\text{base}},\sigma_{\text{disrupt}},\rho_{\text{disrupt}}\}$ were selected using penalized MML. Full details on penalization, kernel, and mean functions are provided in Supplementary Section 6.
@@
Posterior estimates for the cumulative function $F$ (long-term trend plus starvation effect) are shown in Figure 3. Posterior estimates for the decomposed starvation effect are shown in Figure 4. Consistent with prior reports, the Rikenellaceae family demonstrates a substantial decrease during the initial starvation period, followed by a slow recovery to baseline (Silverman et al. 2018, Sun & Zhou 2024). This finding is notable given that decreases in Rikenellaceae are also observed in obese individuals (Okeke et al. 2014, Peters et al. 2018, Tavella et al. 2021). As obesity and starvation seem like opposite conditions, decreased Rikenellaceae relative abundance in both conditions suggests that host factors not present in these artificial gut systems but present in in vivo obesity studies (e.g., the host immune system) may be driving the decreases in obesity.
@@
Our MultiAddGPs also reveal new features of this data and suggest novel biology. For example, our model finds strong evidence that the Rikenellaceae relative abundance in the community decreases drastically during the initial starvation event, followed by a later over-correction where their abundance ultimately increases before trending back to pre-starvation levels. These types of over-corrections are predicted by ecological predator-prey models, which can lead to damped oscillations in response to exogenous stimuli (Samuelson 1971). To best of our knowledge, this is the first study to demonstrate phenomena consistent with that theory within in vitro systems.
@@
Figure 3: MultiAddGPs Posterior for the Smoothed State $\mathbf{F}$ of an Artificial Gut Study. As described in the main text, $\mathbf{F}$ was modeled as the additive combination of a slowly-varying temporal trend and a more variable yet short-lived starvation effect in Vessels 1 and 2 which started on Day 11. Posterior mean and 95% credible intervals of $\mathbf{F}$ are shown for each of the four artificial gut vessels. We plot these results with respect to the Centered Log-Ratio (CLR) Coordinates of each bacterial family. The starvation event, which occurred between Days 11 and 13 in Vessels 1 and 2, is highlighted in Gray.
@@
Figure 4: MultiAddGPs Posterior for the Starvation effect ($\mathbf{f}^{\text{(disrupt)}}$) in Artificial Gut Study. Posterior mean, 50% and 95% credible intervals are shown for all taxa.
5 Application to Longitudinal Microbiome Data
We reanalyzed a human gut microbiome study wherein the study participant collected daily fecal samples over the course of one year (David, Materna, Friedman, Campos-Baptista, Blackburn, Perrotta, Erdman & Alm 2014). While the original dataset had samples from two donors, we focus on the gut microbiome data from Subject A, as this subject demonstrated a substantial shift in community composition during a period of prolonged international travel (David, Materna, Friedman, Campos-Baptista, Blackburn, Perrotta, Erdman & Alm 2014). While this study has been analyzed by multiple authors (Gibbons et al. 2017, Duvallet et al. 2017), we are the first to produce direct estimates of the effect of travel, partitioning those effects from background temporal variation (see Figure 5).
Data was obtained from European Bioinformatics Institute (EBI) with accession number SAMEG179160 and processed following the methods of David, Materna, Friedman, Campos-Baptista, Blackburn, Perrotta, Erdman & Alm (2014), with the exception that we used DADA2 (Callahan et al. 2016) to call sequence variants rather than the OTU calling methods used in the original study. In total, we analyzed 47 bacterial families over 263 time-points. We modeled these data as consisting of two overlapping temporal components: a long-term trend, denoted as $\mathbf{f}^{(\text{trend)}}$, and a non-stationary function representing the effect of travel intervention, denoted as $\mathbf{f^{(\text{travel})}}$. The data were analyzed using the following MultiAddGPs model:
$\mathbf{Y}_{\cdot n}$ $\sim\text{Multinomial}(\boldsymbol{\Pi}_{\cdot n})$
$\boldsymbol{\Pi}_{\cdot n}$ $=\text{ALR}_{D}^{-1}(\mathbf{H}_{\cdot n})$
$\mathbf{H}_{\cdot n}$ $\sim N(\mathbf{F}_{\cdot n},\boldsymbol{\Sigma})$
$\mathbf{F}$ $=\beta_{0}+\mathbf{f}^{(\text{trend})}(t_{n})+\mathbf{f}^{(\text{travel})}(t_{n})$
$\mathbf{f}^{(\text{trend})}$ $\sim\text{GP}(\boldsymbol{\Theta}^{(\text{trend})},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{(\text{trend})})$
$\mathbf{f}^{(\text{travel})}$ $\sim\text{GP}(\boldsymbol{\Theta}^{(\text{travel},v)},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{(\text{travel})})$
$\boldsymbol{\Sigma}$ $\sim\text{InvWishart}(\boldsymbol{\Xi},\zeta)$
where $\boldsymbol{\Xi}$ was chosen following Silverman et al. (2022) to reflect the phylogenetic relationships between taxa. $\zeta=D+100$ was chosen to reflect moderately strong prior knowledge of the importance of these relationships. A squared exponential kernel ($\mathbf{\Gamma}^{\text{trend}}$) was used to capture long-term nonlinear trends, while a non-stationary kernel with a non-linear warp function ($\mathbf{\Gamma}^{\text{travel}}$) used to capture the shorter-term non-linear effect of travel. Hyperparameters $\Omega=\{\sigma_{\text{trend}},\rho_{\text{trend}},\sigma_{\text{travel}},\rho_{\text{travel}}\}$ were estimated using penalized MML. Figure 5 shows the contribution of each component and the cumulative effects. Consistent with David, Maurice, Carmody, Gootenberg, Button, Wolfe, Ling, Devlin, Varma, Fischbach et al. (2014), our findings indicate that while the microbiome remains relatively stable over the long term, travel events induce significant perturbations in community composition. Full details of the analysis are provided in Supplementary Section 7.
Figure 5: Posterior of each component and their cumulative effect for bacterial families Fusobacteriaceae, Monoglobaceae, and Yersiniaceaethe (Longitudinal microbiome data). Each row shows contribution of individual components ($\mathbf{f}^{\text{(trend)}}$ and $\mathbf{f}^{\text{(travel)}}$) and the final column presents the cumulative effect ($\mathbf{f}^{\text{(trend)}}+\mathbf{f}^{\text{(travel)}}$). Results are reported in terms of Centered Log-Ratio (CLR) coordinates, along with posterior means and corresponding 50% as well as 95% credible intervals.